OpenStudy (anonymous):

how do i do problems like (y+5)(y+2)+0?

OpenStudy (anonymous):

I think you mean =0. And recall that if (y+a)(y+b) = 0 then either y+a = 0 or y+b = 0.

OpenStudy (anonymous):

yes, i meant =0 but i still have no idea what ur saying!

OpenStudy (anonymous):

i mean..i know what you mean, but that doesn't help really

OpenStudy (anonymous):

well, if y+5 = 0 what is y ?

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

Yep. Now if y+2 = 0 what is y?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Right, so the solutions to (y+5)(y+2) = 0 are y = -5 and y = -2

OpenStudy (anonymous):

If y is either of those values, the expression (y+5)(y+2) will equal 0.

OpenStudy (anonymous):

dont you have to foil or whatever?

OpenStudy (anonymous):

No, you can solve by inspection.

OpenStudy (anonymous):

i just typed this huge long question and then checked your answer in the back of my math book, and you were right! thanks!